Between 1 and 31 are inserted in AM's so  that the ratio of the 7th and( n -1)the means is 5:9. then the value of the n

Grade:11

2 Answers

Arun
5 years ago
Dear student
 
31 = 1 + (n+1) d
 
(n+ 1) d = 30
 
Now
 
Also
 
1 + 7d /1 + (n-1) d = 5/9
 
Now there rae two equation and two variables.now you can solve it.
Arun
5 years ago
Dear students
 
31 = 1 + (n+1) d
30 = (n+1) d
 
also
 
1 + 7 d / 1+ (n- 1) d = 5/9
 
now there are 2 equations and 2 variables.